OpenStudy (anonymous):

Please display the steps of how you solved a and b In an experiment, a petri dish with a colony of bacteria is exposed to cold temperatures and then warmed again. Time (hours) 0 1 2 3 4 5 6 Population (1000s) 5.1 3.03 1.72 1.17 1.38 2.35 4.08 a. Find a quadratic model for the data in the table. Type your answer below. o b. Use the model to estimate the population of bacteria at 9 hours. Type your answer below.

OpenStudy (whpalmer4):

General form for a quadratic is \[y = ax^2 +bx + c\]although in this case we might do \[p = at^2+bt+c\] We need to find values for \(a,b,c\) to fit our data table. We'll plug in select values from the table to create a system of equations we can solve to find those values. The most obvious value to plug in is \(t=0\): \[5.1 = a(0)^2+b(0)+c\]That gives us a value for \(c\). Now plug in \(t = 1\) and our value for \(c\): \[3.03 = a(1)^2+b(1) + c\] That gives us an equation in terms of \(a\) and \(b\) (remember that we've replaced \(c\) with the numeric value we found in the first step) Repeat with another point in the table to get a second equation in terms of \(a\) and \(b\). Solve that system of two equations for the values of \(a\) and \(b\), write out your completed equation, and use it to predict the population at \(t = 9\).

OpenStudy (anonymous):

I don't know what's going on for a but for B do I basically just plug in 9 as x or t to get the estimated population?

OpenStudy (whpalmer4):

Let's work through it, then. at t = 0, the population is 5.1, so \[5.1 = at^2+bt+c\]\[5.1=a(0)^2+b(0)+c\]\[5.1=0+0+c\]\[5.1=c\]Agreed?

OpenStudy (anonymous):

Agreed

OpenStudy (whpalmer4):

That makes our new, improved formula \[p(t) = at^2+bt+5.1\]Let's plug in another value, how about t=1, p(1) = 3.03: \[3.03 = a(1)^2+b(1)+5.1\]\[3.03= a+b+5.1\] Set that equation aside for the moment, but we'll use it again shortly. Plug in yet another value, how about t = 2, p(2) = 1.72 \[1.72 = a(2)^2+b(2) + 5.1\]\[1.72=4a+2b+5.1\] Now we have a system of two equations in two unknowns to solve: \[3.03 = a + b + 5.1\]\[1.72=4a+2b+5.1\]I would suggest multiplying the first equation by -2, and adding them together. Then you'll have an equation in only one variable, which is easily solved.

OpenStudy (whpalmer4):

Do that and report what you get...

OpenStudy (anonymous):

>.< I don't know what exactly I'm multiplying... I mean you said to multiply the first equation by -2 but I don't see where that... Am I multiplaying like this? 3.03(-2) = a(-2) + b(-2) + 5.1(-2) Because I really don't see any other way to do it...

OpenStudy (whpalmer4):

Yes, that's exactly what I meant. We're solving by elimination. You could also solve the first equation for \(a\), and then substitute what you got for \(a\) in the other equation, if that's more familiar (seems like more work to me, but I'm not the one doing it!) If you went ahead as I suggested, you would get: \[-6.06 -2a-2b-10.2\]\[1.72=4a+2b+5.1\]Adding them together \[1.72-6.06 = 4a-2a+2b-2b+5.1-10.2\]\[1.72-6.06=2a+5.1-10.2\]and I know you can solve that. Take the value you get for \(a\) and plug it back into either of the equations (I would suggest \(3.03 = a + b + 5.1\) is the easiest to use) to find the value of \(b\). Write out \(p(t) = at^2+bt+c\) with your values of \(a,b,c\) in place of the letters and you're done with the first part. For the second part, you just evaluate \(p(9) = a(9)^2+b(9) +c\) with your values of \(a,b,c\).